OpenStudy (anonymous):

write an explicit definition for the sequence -2,1,4,7,...

OpenStudy (anonymous):

on subtracting 2nd term (1) with 1st term(-2) I got 3 similarly on subtracting 3rd term (4) with 2nd term (1) I got again 3 Which suggest its an Arithmetic series as the common difference is same so general equation of this series is: \[\alpha= a+(n-1)*d where, \alpha is the nth term and a is 1st term, is the common difference\]

OpenStudy (anonymous):

is it an=2+(n=3)

OpenStudy (anonymous):

or 2=(n-3)

OpenStudy (anonymous):

2+(n-3)

OpenStudy (anonymous):

I met -2

OpenStudy (anonymous):

yes its right @bowershe

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

the choices are an=-2+3n an=-2+(n-3) an=-2+(n+3) an=2+3(n-1)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

ok